How far shall an object fall for it to have a velocity of 3.50 m/s? Neglect air resistance. 




1
Expert's answer
2021-11-16T19:20:19-0500

"L=\frac{V^2}{2g}" ; V=3.50m/s

"L=\frac{3.50^2}{2\times 9.8}=0.625m"